a truck driver drives from city A to city B in 5hr at a certain average speed and drives back from city B to city A in 4 hr at an average speed of 15 miles per faster, which of the following is the truck driver average speed from city A to city B
Answer by ankor@dixie-net.com(22740)

A truck driver drives from city A to city B in 5 hr at a certain average speed and drives back from city B to city A in 4 hr at an average speed of 15 miles per faster, what is the truck driver average speed from city A to city B
:
let s = his speed from A to B
then
(s+15) = his speed from B to A
:
The distance is the same each way, write a dist equation dist = time * speed
5s = 4(s+15)
5s = 4s + 60
5s - 4s = 60
s = 60 mph is his speed from A to B
:
;
Confirm this, find the actual distance (return speed 60+15)
5 * 60 = 300 mi
4 * 75 = 300 mI